12. (4) Graph the system of inequalities.
x – 4y < 0
x + y > 0



Answer by pvaka(23)

x-4y<0 GRAPH FORM y>(1/4)x (m=1/4)
x+y>0 GRAPH FORM y>-x (m=-1)